OpenStudy (anonymous):

y=3sinx+sin^2(x) find the maximum and minimum values

OpenStudy (anonymous):

apparently the max is 4 and the min in -2 maybe we can take the derivative and find the critical points

OpenStudy (anonymous):

actually that is unnecessary forget i mentioned it

OpenStudy (anonymous):

biggest \[3\sin(x)\] can be is 3, since \[-1\leq\sin(x)\leq 1\] and similarly the smallest it can be is -3

OpenStudy (anonymous):

likewise the largest \[\sin^2(x)\] can be is also 1

OpenStudy (anonymous):

max/min values occur at x = pi/2 and 3pi/2. To find the actual values, stick these x values into the original equation.

OpenStudy (anonymous):

so largest \[3\sin(x)+\sin^2(x)\] can be is 4

OpenStudy (anonymous):

minimum occurs when \[\sin(x)=-1\] meaning \[3\sin(x)=-3\] and \[sin^2(x)=1\] so minimum is -2

OpenStudy (anonymous):

you can use calculus if you like, but thinking also works for this question, no calc needed

OpenStudy (earthcitizen):

With calculus y=3sinx+sin^2(x) at stationary point y'=0 y'=2 sin(x)+3) cos(x)=0

OpenStudy (anonymous):

the derivative is y'=3cosx+2sinxcosx

OpenStudy (anonymous):

don't forget the chain rule...

OpenStudy (earthcitizen):

yh, my bad y'=(2 sin(x)+3) cos(x)

OpenStudy (earthcitizen):

y'=(2 sin(x)+3) cos(x)=0